the length of a rectangle is three more than twice the width determine the dimensions that will give a total area of 27m^2
please help,i'm completely stuck

Answer by lwsshak3(11628)

the length of a rectangle is three more than twice the width determine the dimensions that will give a total area of 27m^2
***
let x=width of rectangle
2x+3=length of rectangle
area=length*width
..
x(2x+3)=27
2x^2+3x=27
2x^2+3x-27=0
(2x+9(x-3)=0
x=-9/2 (reject, x>0)
or
x=3
2x+3=9
width of rectangle=3m
length of rectangle=9m